OpenStudy (anonymous):

solutions again ... y-5x=-6 3y-15x=-12

hero (hero):

Hint: Divide both sides of the second equation by 3

OpenStudy (anonymous):

how? i dont get what u mean

hero (hero):

Really bro?

OpenStudy (anonymous):

yes. i am bad at this crap

hero (hero):

3(y - 5x) = 3(-4) y - 5x = -4

hero (hero):

Now do you see?

OpenStudy (anonymous):

i think so

hero (hero):

So you're left with y - 5x = -6 y - 5x = -4 Do you know what this means?

OpenStudy (anonymous):

no solutions?

hero (hero):

Exactly

OpenStudy (anonymous):

thanks!

hero (hero):

Yeah, I do that first to make sure there are no solutions. Usually if you are able to reduce one or both equations to a simpler form, chances are you'll either have no solutions or at the very least, solving the system will be easier.

OpenStudy (anonymous):

ok thank you